Hi, my daughter is in 7th grade PreAlgebra and we do not have our book yet. She has been given her assignments as typed handouts. One particular problem has both of us STUMPED. The teacher's instructions were to use an equation to answer this problem. Thanks for your help.
3 bacteria are placed in a petri dish and then counted each hour. The first three bacteria counts were: 12, 39, 120. If this pattern continues, how many hours after the bacteria are placed in the dish will there be 3279 bacteria?
(A) after 6 hours (B) after 12 hours (C) after 8 hours (D) after 4 hours

Answer by Nate(3500)

You are taking the original times three plus three:
3x + 3
0 hours ~> 3
1 hour ~> 3*3 + 3 = 12
2 hours ~> 3*12 + 3 = 39
3 hours ~> 3*39 + 3 = 120
4 hours ~> 3*120 + 3 = 363
5 hours ~> 3*363 + 3 = 1092
6 hours ~> 3*1092 + 3 = 3279
Answer: (A)